# Summarization Metrics for Spanish and Basque: Do Automatic Scores and LLM-Judges Correlate with Humans?

### v1

@@ Abstract
Studies on evaluation metrics and LLM-as-a-Judge models for automatic text summarization have largely been focused on English, limiting our understanding of their effectiveness in other languages. Through our new dataset BASSE (BAsque and Spanish Summarization Evaluation), we address this situation by collecting human judgments on 2,040 abstractive summaries in Basque and Spanish, generated either manually or by five LLMs with four different prompts. For each summary, annotators evaluated five criteria on a 5-point Likert scale: coherence, consistency, fluency, relevance, and 5W1H. We use these data to reevaluate traditional automatic metrics used for evaluating summaries, as well as several LLM-as-a-Judge models that show strong performance on this task in English. Our results show that currently proprietary judge LLMs have the highest correlation with human judgments, followed by criteria-specific automatic metrics, while open-sourced judge LLMs perform poorly. We release BASSE [^1] The dataset is available under open licenses at https://github.com/hitz-zentroa/summarization. and our code publicly, along with the first large-scale Basque summarization dataset containing 22,525 news articles with their subheads.
@@ 1 Introduction
Automatic text summarization seeks to create a concise and fluent text that captures the main ideas from one or more documents (?). This task has been studied in depth in Natural Language Processing (NLP) (?, ?) and plays an important role in mitigating excess information in an age of increasing information glut (?).
@@
Although extractive approaches, where sentences and phrases are taken directly from the original document, were originally more common (?, ?), current approaches instead generally rely on abstractive methods, where the model generates novel sentences to capture the most salient information. This change in modeling paradigm also leads to new challenges for evaluation, as abstractive models are more prone to hallucinating information (?, ?, ?).
@@
In order to evaluate automatic summaries, researchers generally rely on three main strategies: i) human evaluation, ii) evaluation using reference summaries and automatic metrics, and iii) evaluation via Large Language Models (LLMs), often referred to as LLM-as-a-Judge. Human evaluation over a small number of important criteria, e.g., coherence or fluency, is often chosen to give high-quality evaluation of abstractive summaries. However, this approach is slow, expensive, and does not lend itself to internal evaluations while iterating over models. Therefore, most research instead relies on automatic metrics, e.g., ROUGE (?) and, more recently, LLM-as-a-Judge (?, ?, ?).
@@
Selecting appropriate metrics or judge LLMs relies on meta-evaluations (?, ?, ?) that measure the correlation between human evaluations and automatic evaluation methods. While these studies provide invaluable information, they have only been performed comprehensively for English, leaving the assessment of their performance in other languages to be explored.
@@
We therefore address this gap by collecting human annotations for automatic summaries in Basque and Spanish—languages with respective fragmentary and medium coverage of language resources (?). Specifically, we collect abstractive summaries generated by five LLMs and four different prompts on 45 news articles in each language. We then collect human judgments for each summary with respect to five criteria—namely, coherence, consistency, fluency, relevance, and 5W1H. With this novel dataset, BAsque and Spanish Summarization Evaluation (BASSE), we then proceed to address the following research questions:
  1. RQ1: Do automatic metrics for evaluating summarization correlate well with human judgments in Basque and Spanish?
  2. RQ2: Do judge LLMs have the ability to replicate human judgments for summarization?
@@
  3. RQ3: Which summary criteria (coherence, consistency, fluency, relevance, and 5W1H) do state-of-the-art LLMs struggle with in Basque and Spanish?
@@
Our results indicate that proprietary LLM-as-a-Judge models currently have the highest correlation with human judgments on our data, while several traditional metrics also correlate well with specific criteria. On the contrary, open-source judge LLMs currently do not correlate well with human judgments on our data. Our contributions are hence the following:
@@ 2 Related Work
Two main approaches exist in summarization: extractive and abstractive (?). The extractive approach selects and takes the most relevant sentences directly from the input document, concatenating them to create the summary. On the other hand, the abstractive approach generates a set of novel sentences to represent the main information of the original document, rather than copying directly from the original text. In the case of automatic summarization, abstractive summaries are “more similar to those produced by humans” (?) and have become more popular as summarization models have improved.
English datasets for training and evaluating summarization models range from the original Document Understanding Conference (DUC) and Text Analysis Conference (TAC) shared tasks (?, ?), which are relatively small but high quality, to automatically collected datasets, such as CNN/DailyMail (?) or XSUM (?), which rely on collecting news articles paired with their subhead as a proxy for a human-generated summary.
Non-English languages have fewer resources available and most use news articles paired with their subhead (?, ?). There are notable exceptions that instead collect human-generated summaries (?, ?), but for most languages there are no high-quality resources available for evaluating automatic summaries.
@@ 2.1 Manual Evaluation
Manual evaluation of automatic summaries has long been the gold standard approach, ranging from direct annotation of criteria to the pyramid method (?). This evaluation can be done through directly annotating each summary for certain criteria—coherence of the generated text, relevance of the selected material, etc.—or alternatively by ranking automatic summaries with respect to one of these criteria. In addition, as abstractive summaries have become more common, evaluating their consistency (factuality/faithfulness) has become increasingly important as well (?, ?, ?).
Importantly, crowdsourcing often leads to poor results when evaluating automatic summarization (?, ?, ?) and expert evaluation can also fail to achieve good levels of inter-annotator agreement (?, ?). This can be mitigated to a degree by setting up well-defined annotation guidelines in order to achieve reliable human evaluation of text summarization.
@@ 2.2 Automatic Evaluation
Manual evaluation becomes an impossible task as corpus size increases, motivating the need for automatic metrics for summarization evaluation. Reference-based metrics include ROUGE (?) and its many variants (?, ?), which use n-gram overlaps between reference and automatic summaries. S3 (?) is instead a learned metric that combines the output of other metrics as features and learns to predict human scores directly. Similarly, referenced-based Natural Language Generation (NLG) evaluation metrics, such as BLEU (?), BertScore (?), CHRF, or CIDEr, are also commonly used.
Reference-free metrics have also been proposed (?, ?, ?). SummaQA (?), for example, automatically creates questions from the source document and scores the summary depending on its ability to answer these questions. BLANC (?) instead measures how much a summary helps a language model predict words from a sentence taken from the original document.
More recently, using LLMs to automatically evaluate summaries—LLM-as-a-Judge—has shown promise (?, ?, ?). These models often show higher correlations with human judgments than previous token-based metrics.
However, nearly all of these metrics and models have been designed for English and their ability to correctly evaluate non-English summaries has not been deeply studied. In this line, ? (?) highlight the importance of not relying solely on summarization evaluation metrics conceived for English texts. It is furthermore nontrivial to transfer many of these metrics to new languages, as they require a number of resources to already exist.
@@ 2.3 Meta-evaluation of Automatic Metrics
Meta-evaluation of automatic metrics is necessary to establish the reliability of the metrics. This is typically done by studying the correlation between metric scores and human judgments (?).
? (?), for example, collect human evaluations for 23 English summarization models over four dimensions: coherence, consistency, fluency, and relevance. They then calculate correlations between these annotations and automatic metrics and find that few automatic metrics correlate well with the human annotations.
Despite some studies on Spanish and French (?), there has been little meta-evaluation done on other languages. ? (?) create a multilingual summary dataset with six languages and also collect binary yes/no crowd-sourced annotations for six criteria, although interannotator agreement (IAA) is quite low, with Krippendorf’s $\alpha$ below 0.5 for 5 of 6 criteria. They then train a referenceless metric and compare it with ROUGE-L, but do not perform a comprehensive study of evaluation metrics. Similarly, ? (?) create a multilingual summary dataset of 45 languages and crowd-source annotations for two criteria. They then compare ROUGE-L and an NLI-based faithfulness metrics on these data. Finally, ? (?) meta-evaluate a number of automatic metrics on summaries generated by BERT and pointer-network models in eight languages.
However, none of the meta-evaluations performed on non-English languages study a range of automatic metrics and most of them use automatically extracted subhead-article pairs as a proxy for reference summaries, which limits the reliability of the findings.
@@ 3.1 Data Collection
For Basque, we collect 45 articles from Berria, [^2] https://www.berria.eus the main newspaper in this language, including metadata concerning the URL, date of publication, assigned topic category, title of the article, subhead, as well as the body of the text. [^3] The documents were processed with trafilatura (?). Berria offers its contents under CC BY-SA 4.0 license. [^4] In addition to the 45 articles we annotate, we also collect 22,525 Basque articles from Berria with their subhead which we release with BASSE.
For Spanish, another set of 45 news articles was randomly extracted from the test set of the MLSUM dataset (?). The Spanish subset of MLSUM contains news articles of El País newspaper [^5] https://www.elpais.com and provides (news body, subhead, URL) triplets. MLSUM is distributed under MIT license, complying with our needs for redistribution.
@@ 3.2 Summary Collection
  • Subhead: Following standard practice in news summarization literature (?, ?), we extracted the original subheads from each article to serve as baseline summaries.
@@ 3.2.1 Models
We selected a number of proprietary and open-source LLMs that are able to provide summaries in Basque and Spanish. Proprietary models included Anthropic’s Claude, [^6] https://www.anthropic.com/claude OpenAI’s GPT-4o, [^7] https://openai.com/index/gpt-4o-system-card and Reka AI’s Reka (?). [^8] We used Claude 3.5 Sonnet for Rounds 1, 2, and 3 and Claude 3.5 Haiku for Round 4. Similarly, we used Reka-Core for Rounds 1,2, and 3 and Reka-Flash for Round 4. The open-source options were Cohere’s Command R+ [^9] https://docs.cohere.com/v2/docs/command-r-plus and Meta’s Llama-3.1-70b-Instruct (?). All models were prompted using their default configurations through their respective native chat interfaces, with the exception of Llama, which was accessed through HuggingFace’s chat interface [^10] https://huggingface.co/chat due to the absence of a proprietary chat platform.
@@ 3.2.2 Prompts
Along with the selection of those five different models, we experimented with the following four prompts to collect summaries from the LLMs. These prompts, presented below, were designed in order to obtain fairly diverse summaries. Although multilingual LLMs often show stronger performance when prompted in English (?, ?), in early experiments, we found that prompting the LLMs in English led to the summaries also being given in English. Therefore, we opted to prompt the model directly in the language in which it should write the summary. Table 1 shows each prompt in Basque and Spanish, as well as its translation to English and their descriptions are provided below.
@@
  • 5W1H: This prompt, which is a further derivation of Core, uses an approach common in journalism consisting of answering six key questions: the five “W” questions (who, what, when, where, why) and the 1 “H” question (how). This technique “clearly describe[s] key information of news in an explicit manner” (?) and allows to capture “the semantic constituents of a sentence which are comparatively simpler to understand and identify” (?). The model should then use this information to construct the summary for the following document.
@@ 3.3 Human Annotation
We followed the annotation procedure of ? (?), who based their criteria largely on ? (?) and ? (?). For our experiment, besides the following descriptions of the used criteria, we elaborated a detailed set of annotation guidelines, which we provide in Annex A. In addition, Annex B contains examples of annotated summaries.
@@ 3.3.1 Criteria
While maintaining the core criteria of previous work (?), we extended the framework with an additional criterion, 5W1H, which we considered necessary to determine whether the summary contains all relevant information from the source document. The five criteria used in our evaluation are defined as follows:
@@
  • Fluency scores the quality of individual sentences. Drawing again from the DUC quality guidelines, sentences in the summary “should have no formatting problems, capitalization errors or obviously ungrammatical sentences (e.g., fragments, missing components) that make the text difficult to read” (?).
@@
  • 5W1H measures how well important information is preserved from the source by following a journalistic approach. It evaluates whether key content is presented completely by addressing the six fundamental questions (see prompt 5W1H in Section 3.2.2). The summary should not lack any important information available in the source document. This recall-oriented criterion corresponds roughly to coverage in ? (?), although it differs in that we calculate 5W1H relative to the original document, whereas they use the reference summary to calculate coverage.

### v2

@@ Abstract
Studies on evaluation metrics and LLM-as-a-Judge models for automatic text summarization have largely been focused on English, limiting our understanding of their effectiveness in other languages. Through our new dataset BASSE (BAsque and Spanish Summarization Evaluation), we address this situation by collecting human judgments on 2,040 abstractive summaries in Basque and Spanish, generated either manually or by five LLMs with four different prompts. For each summary, annotators evaluated five criteria on a 5-point Likert scale: coherence, consistency, fluency, relevance, and 5W1H. We use these data to reevaluate traditional automatic metrics used for evaluating summaries, as well as several LLM-as-a-Judge models that show strong performance on this task in English. Our results show that currently proprietary judge LLMs have the highest correlation with human judgments, followed by criteria-specific automatic metrics, while open-sourced judge LLMs perform poorly. We release BASSE [^1] BASSE is available under open licenses at https://github.com/hitz-zentroa/summarization. and our code publicly, along with the first large-scale Basque summarization dataset containing 22,525 news articles with their subheads.
@@ 1 Introduction
Automatic text summarization seeks to create a concise and fluent text that captures the main ideas from one or more documents Giarelis, Mastrokostas, and Karacapilidis (2023). This task has been studied in depth in Natural Language Processing (NLP) Dang (2006); Dang and Owczarzak (2008) and plays an important role in mitigating excess information in an age of increasing information glut Navas-Martin, Albornos-Muñoz, and Escandell-García (2012).
@@
Although extractive approaches, where sentences and phrases are taken directly from the original document, were originally more common McDonald (2007); Lin and Bilmes (2011), current approaches instead generally rely on abstractive methods, where the model generates novel sentences to capture the most salient information. This change in modeling paradigm also leads to new challenges for evaluation, as abstractive models are more prone to hallucinating information Maynez et al. (2020); Vasilyev, Dharnidharka, and Bohannon (2020); Laban et al. (2022).
@@
In order to evaluate automatic summaries, researchers generally rely on three main strategies: i) human evaluation, ii) evaluation using reference summaries and automatic metrics, and iii) evaluation via Large Language Models (LLMs), often referred to as LLM-as-a-Judge. Human evaluation over a small number of important criteria, e.g., coherence or fluency, is often chosen to give high-quality evaluation of abstractive summaries. However, this approach is slow, expensive, and does not lend itself to internal evaluations while iterating over models. Therefore, most research instead relies on automatic metrics, e.g., ROUGE Lin (2004) and, more recently, LLM-as-a-Judge Leiter et al. (2023); Kim et al. (2024); Lyu et al. (2025).
@@
Selecting appropriate metrics or judge LLMs relies on meta-evaluations Bhandari et al. (2020); Fabbri et al. (2021); Deutsch, Dror, and Roth (2022) that measure the correlation between human evaluations and automatic evaluation methods. While these studies provide invaluable information, they have only been performed comprehensively for English, leaving the assessment of their performance in other languages to be explored.
@@
We therefore address this gap by collecting human annotations for automatic summaries in Basque and Spanish—languages with respective fragmentary and medium coverage of language resources Giagkou et al. (2023). Specifically, we collect abstractive summaries generated by five LLMs and four different prompts on 45 news articles in each language. We then collect human judgments for each summary with respect to five criteria—namely, coherence, consistency, fluency, relevance, and 5W1H. With this novel dataset, BAsque and Spanish Summarization Evaluation (BASSE), we then proceed to address the following research questions:
@@
  • RQ3: Which summary criteria (coherence, consistency, fluency, relevance, and 5W1H) do state-of-the-art LLMs struggle with in Basque and Spanish?
@@
Our results indicate that proprietary LLM-as-a-Judge models currently have the highest correlation with human judgments on our data, while several traditional metrics also correlate well with specific criteria. On the contrary, open-source judge LLMs currently do not correlate well with human judgments on our data.
Our contributions are hence the following:
@@ 2 Related Work
Two main approaches exist in summarization: extractive and abstractive El-Kassas et al. (2021). The extractive approach selects and takes the most relevant sentences directly from the input document, concatenating them to create the summary. On the other hand, the abstractive approach generates a set of novel sentences to represent the main information of the original document, rather than copying directly from the original text. In the case of automatic summarization, abstractive summaries are “more similar to those produced by humans” Giarelis, Mastrokostas, and Karacapilidis (2023) and have become more popular as summarization models have improved.
English datasets for training and evaluating summarization models range from the original Document Understanding Conference (DUC) and Text Analysis Conference (TAC) shared tasks Dang (2006); Dang and Owczarzak (2008), which are relatively small but high quality, to automatically collected datasets, such as CNN/DailyMail Hermann et al. (2015) or XSUM Narayan, Cohen, and Lapata (2018), which rely on collecting news articles paired with their subhead as a proxy for a human-generated summary.
Non-English languages have fewer resources available and most use news articles paired with their subhead Scialom et al. (2020); Varab and Schluter (2021). There are notable exceptions that instead collect human-generated summaries Segarra Soriano et al. (2022); Gonzalez-Agirre et al. (2024), but for most languages there are no high-quality resources available for evaluating automatic summaries.
@@ 2.1 Manual Evaluation
Manual evaluation of automatic summaries has long been the gold standard approach, ranging from direct annotation of criteria to the pyramid method Nenkova and Passonneau (2004). This evaluation can be done through directly annotating each summary for certain criteria—coherence of the generated text, relevance of the selected material, etc.—or alternatively by ranking automatic summaries with respect to one of these criteria. In addition, as abstractive summaries have become more common, evaluating their consistency (factuality/faithfulness) has become increasingly important as well Maynez et al. (2020); Vasilyev, Dharnidharka, and Bohannon (2020); Laban et al. (2022).
Importantly, crowdsourcing often leads to poor results when evaluating automatic summarization Gillick and Liu (2010); Hardy, Narayan, and Vlachos (2019); Fabbri et al. (2021) and expert evaluation can also fail to achieve good levels of inter-annotator agreement Steen and Markert (2021); Iskender, Polzehl, and Möller (2021). This can be mitigated to a degree by setting up well-defined annotation guidelines in order to achieve reliable human evaluation of text summarization.
@@ 2.2 Automatic Evaluation
Manual evaluation becomes an impossible task as corpus size increases, motivating the need for automatic metrics for summarization evaluation. Reference-based metrics include ROUGE Lin (2004) and its many variants Rankel et al. (2013); Graham (2015), which use n-gram overlaps between reference and automatic summaries. S3 Peyrard, Botschen, and Gurevych (2017) is instead a learned metric that combines the output of other metrics as features and learns to predict human scores directly. Similarly, referenced-based Natural Language Generation (NLG) evaluation metrics, such as BLEU Papineni et al. (2002), BertScore Zhang et al. (2020), CHRF, or CIDEr, are also commonly used.
Reference-free metrics have also been proposed Louis and Nenkova (2009a, b); Saggion et al. (2010). SummaQA Scialom et al. (2019), for example, automatically creates questions from the source document and scores the summary depending on its ability to answer these questions. BLANC Vasilyev, Dharnidharka, and Bohannon (2020) instead measures how much a summary helps a language model predict words from a sentence taken from the original document.
More recently, using LLMs to automatically evaluate summaries—LLM-as-a-Judge—has shown promise Leiter et al. (2023); Kim et al. (2024); Lyu et al. (2025). These models often show higher correlations with human judgments than previous token-based metrics.
However, nearly all of these metrics and models have been designed for English and their ability to correctly evaluate non-English summaries has not been deeply studied. In this line, Lloret, Plaza, and Aker (2018) highlight the importance of not relying solely on summarization evaluation metrics conceived for English texts. It is furthermore nontrivial to transfer many of these metrics to new languages, as they require a number of resources to already exist.
@@ 2.3 Meta-evaluation of Automatic Metrics
Meta-evaluation of automatic metrics is necessary to establish the reliability of the metrics. This is typically done by studying the correlation between metric scores and human judgments Lin (2004).
Fabbri et al. (2021), for example, collect human evaluations for 23 English summarization models over four dimensions: coherence, consistency, fluency, and relevance. They then calculate correlations between these annotations and automatic metrics and find that few automatic metrics correlate well with the human annotations.
Despite some studies on Spanish and French Saggion et al. (2010), there has been little meta-evaluation done on other languages. Clark et al. (2023) create a multilingual summary dataset with six languages and also collect binary yes/no crowd-sourced annotations for six criteria, although interannotator agreement (IAA) is quite low, with Krippendorf’s $\alpha$ below 0.5 for 5 of 6 criteria. They then train a referenceless metric and compare it with ROUGE-L, but do not perform a comprehensive study of evaluation metrics. Similarly, Aharoni et al. (2023) create a multilingual summary dataset of 45 languages and crowd-source annotations for two criteria. They then compare ROUGE-L and an NLI-based faithfulness metrics on these data. Finally, Koto, Lau, and Baldwin (2021) meta-evaluate a number of automatic metrics on summaries generated by BERT and pointer-network models in eight languages.
However, none of the meta-evaluations performed on languages other than English study a range of automatic metrics and most of them use automatically extracted subhead-article pairs as a proxy for reference summaries, which limits the reliability of the findings.
@@ 3.1 Data Collection
For Basque, we collect 45 articles from Berria, [^2] https://www.berria.eus the main newspaper in this language, including metadata concerning the URL, date of publication, assigned topic category, title of the article, subhead, as well as the body of the text. [^3] The documents were processed with trafilatura Barbaresi (2021). Berria offers its contents under CC BY-SA 4.0 license. [^4] In addition to the 45 articles we annotate, we also collect 22,525 Basque articles from Berria with their subhead which we release with BASSE.
For Spanish, another set of 45 news articles was randomly extracted from the test set of the MLSUM dataset Scialom et al. (2020). The Spanish subset of MLSUM contains news articles of El País newspaper [^5] https://www.elpais.com and provides (news body, subhead, URL) triplets. MLSUM is distributed under MIT license, complying with our needs for redistribution.
@@ 3.2 Summary Collection
  • Subhead: Following standard practice in news summarization literature Hermann et al. (2015); Narayan, Cohen, and Lapata (2018), we extracted the original subheads from each article to serve as baseline summaries.
@@ 3.2.1 Models
We selected a number of proprietary and open-source LLMs that are able to provide summaries in Basque and Spanish. Proprietary models included Anthropic’s Claude, [^6] https://www.anthropic.com/claude OpenAI’s GPT-4o, [^7] https://openai.com/index/gpt-4o-system-card and Reka AI’s Reka Reka Team et al. (2024). [^8] We used Claude 3.5 Sonnet for Rounds 1, 2, and 3 and Claude 3.5 Haiku for Round 4. Similarly, we used Reka-Core for Rounds 1,2, and 3 and Reka-Flash for Round 4. The open-source options were Cohere’s Command R+ [^9] https://docs.cohere.com/v2/docs/command-r-plus and Meta’s Llama 3.1 70B Instruct Dubey et al. (2024). All models were prompted using their default configurations through their respective native chat interfaces, with the exception of Llama, which was accessed via HuggingFace’s chat interface [^10] https://huggingface.co/chat due to the absence of a proprietary chat platform.
@@ 3.2.2 Prompts
Along with the selection of those five different models, we experimented with the following four prompts to collect summaries from the LLMs. These prompts, presented below, were designed in order to obtain fairly diverse summaries. Although multilingual LLMs often show stronger performance when prompted in English Shi et al. (2022); Etxaniz et al. (2024a), in early experiments, we found that prompting the LLMs in English led to the summaries also being given in English. Therefore, we opted to prompt the model directly in the language in which it should write the summary. Table 1 shows each prompt in Basque and Spanish, as well as its translation to English and their descriptions are provided below.
@@
  • 5W1H: This prompt, which is a further derivation of Core, uses an approach common in journalism consisting of answering six key questions: the five “W” questions (who, what, when, where, why) and the 1 “H” question (how). This technique “clearly describe[s] key information of news in an explicit manner” Zhang, Chen, and Ma (2019) and allows to capture “the semantic constituents of a sentence which are comparatively simpler to understand and identify” Chakma et al. (2020). The model should then use this information to construct the summary for the following document.
@@ 3.3 Human Annotation
We followed the annotation procedure of Fabbri et al. (2021), who based their criteria largely on Kryscinski et al. (2019) and Dang (2005). For our experiment, besides the following descriptions of the used criteria, we elaborated a detailed set of annotation guidelines, which we provide in Appendix A. Figure 12 shows real examples of annotated summaries in Basque (examples in Spanish are shown in Appendix B).
Summary: Command R+ with Core prompt {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=red9 Coh & Con Flu Rel 5W1H Avg
2.00 2.67 1.67 2.00 2.00 2.07
Summary: GPT-4o with Base prompt {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=azure9 Coh & Con Flu Rel 5W1H Avg
4.67 4.67 4.67 4.00 5.00 4.60
Figure 1: Examples of low- and high-quality summaries of a news article in Basque. Scores are the average of three annotators. Source of the article: Berria. [^12] https://www.berria.eus/euskal-herria/ehunka-herritarrek-etxebarriko-sexu-erasoa-salatu-dute_2126343_102.html
@@ 3.3.1 Criteria
While maintaining the core criteria of previous work Fabbri et al. (2021), we extended the framework with an additional criterion, 5W1H, which we considered necessary to determine whether the summary contains all relevant information from the source document. The five criteria used in our evaluation are defined as follows:
@@
  • Fluency scores the quality of individual sentences. Drawing again from the DUC quality guidelines, sentences in the summary “should have no formatting problems, capitalization errors or obviously ungrammatical sentences (e.g., fragments, missing components) that make the text difficult to read” Dang (2006).
@@
  • 5W1H measures how well important information is preserved from the source by following a journalistic approach. It evaluates whether key content is presented completely by addressing the six fundamental questions (see prompt 5W1H in Section 3.2.2). The summary should not lack any important information available in the source document. This recall-oriented criterion corresponds roughly to coverage in Koto, Lau, and Baldwin (2020), although it differs in that we calculate 5W1H relative to the original document, whereas they use the reference summary to calculate coverage.